OpenStudy (anonymous):

The derivative of \(\sec (x) \equiv \frac{1}{\cos (x)}\) is?

OpenStudy (anonymous):

\[\frac{d}{dx}\sec(x)=\sec(x)\tan(x)\]

OpenStudy (anonymous):

no do it step by step according to this method:- \(f (x)\) has a derivative \(f' (x)\) and \(f (x) \neq 0\) then \(\left( \frac{1}{f} \right)' (x) = \frac{- f' (x)}{\left[ f (x) \right]^2}\)

OpenStudy (anonymous):

\(\frac{? (x)}{\cos^2 (x)}\)

OpenStudy (anonymous):

using the quotient rule? alright. \[u := 1\] \[v := \cos(x)\] \[\frac{d}{dx}\frac{u}{v}=\frac{\frac{du}{dx}v - \frac{dv}{dx}u}{v^2}\] \[\frac{d}{dx}\sec(x)=\frac{d}{dx}\frac{1}{\cos(x)}=\frac{-\sin(x)}{\cos^2(x)}\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

there's a mistake there, it should be \[\frac{\sin(x)}{\cos^2(x)}\]

OpenStudy (anonymous):

ya i know